A 40-year-old man in the U.S. has a 0.242% risk of dying during the next year . An insurance company charges $260 per year for a life-insurance policy that pays a $100,000 death benefit. What is the expected value for the person buying the insurance? Round your answer to the nearest dollar.


1
Expert's answer
2021-10-26T10:31:15-0400

Solution:

Given that the annual charges of the insurance policy is $260.

The total amount of the policy is $100,000.

But he pays yearly $260.

So, the insurance amount after his death will be 100,000 - 260 = $ 99,740

The probability of that person death is 0.242% or 0.00242.

Hence, the probability of that person living is, 1 - 0.00242 = 0.99758

The expected value of the policy "=\Sigma x.P(x)"

"=(99740)(0.00242)+(-260)(0.99758)
\\=\$-18"